Hello can someone please help me with this question:
Rewrite y=2(10)^{t/12} in the form y=a(1+r)^t or y=a(1-r)^t. Round the value of r to the nearest ten-thousandth. Then tell whether the model represents exponential growth or exponential decay.
I have tried, I know that it is exponential growth, but I really don't understand how to rewrite in this form.
Please help, Thanks.

Answer by ikleyn(52785)

.

y = 2*10^(t/12) = 2 * (10^(1/12))^t.


I hope you do understand this identity and do unferstand why it is valid.


Next step  10^(1/12) = 1.2115 = 1 + 0.2115.


Therefore  y = 2*(1+0.2115)^t  is that form you are looking for.

-------------

Solved, answered and explained.